OpenStudy (anonymous):

the formula a= d/g gives the average gas milege a of a vehicle that uses g gallsons of gas to travel d miles. use the formula to find how many gallons of gas a vehicle with average gas mileage of 20.2 miles per gallon will use to travel 75 miles. round your answer to the nearest tenth.

OpenStudy (anonymous):

Just use the values they give you in the formula given... a = d / g a = gas mileage, d = miles, g = gallons

OpenStudy (anonymous):

i dont get it....

OpenStudy (anonymous):

so, you are given that a = 20.2, the distance is d = 75, and you need to find g. 20.2 = 75 / g

OpenStudy (anonymous):

so would you multiply by 75

OpenStudy (anonymous):

You need to solve for g, so you need to not have it in the denominator. Multiply both sides of the equation by "g" 20.2 = 75 / g 20.2 * g = g * 75/g Then simplify the right side: 20.2 * g = 75 Finally, you need to divide both sides by 20.2 to get g all alone on the left: g = 75 / 20.2

OpenStudy (anonymous):

do u need to diveide 75/20.2

OpenStudy (anonymous):

yes, that's right... you divide 75 miles by 20.2 miles per gallon to get the answer.

OpenStudy (anonymous):

so 3.7 gallons?

OpenStudy (anonymous):

That's what I get :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome, glad to help